# MANet: Fine-Tuning Segment Anything Model for Multimodal Remote Sensing Semantic Segmentation

### v1

@@ IV-B Implementation details
All experiments were conducted using PyTorch on a single NVIDIA GeForce RTX 3090 GPU with 24GB of RAM. The stochastic gradient descent algorithm was used to train all the models under consideration with a learning rate of $0.01$, momentum of $0.9$, weight decay of $0.0005$, and a batch size of $10$. The batch size is reduced to $6$ when ViT-H is applied to meet the memory limit. Basic data augmentation techniques, including random rotation and flipping, are applied after sample collection with the sliding window.
To assess the semantic segmentation performance on multimodal remote sensing data, we use overall accuracy (OA), mean F1 score (mF1), and mean intersection over union (mIoU). These standard metrics enable a fair comparison between the proposed MANet and other state-of-the-art methods. Specifically, OA evaluates both foreground classes and the background class, while mF1 and mIoU are calculated for the five foreground classes.
Figure 8: Visualized comparisons on the Potsdam test set with the size of $2000\times 2000$. (a) IRRG images, (b) DSM, (c) Ground Truth, (d) MAResU-Net, (e) vFuseNet, (f) FuseNet, (g) ESANet, (h) SA-GATE, (i) CMGFNet, (j) TransUNet, (k) CMFNet, (l) UNetFormer, (m) MFTransNet, (n) FTransUNet, (o) $\rm{RS^{3}Mamba}$, (p) The proposed MANet. Purple boxes are added to highlight the differences.
@@ IV-C Performance Comparison
We benchmarked the proposed MANet against fourteen state-of-the-art methods, including ABCNet [43], PSPNet [47], MAResU-Net [44], vFuseNet [30], FuseNet [9], ESANet [45], SA-GATE [48], CMGFNet [46], TransUNet [49], CMFNet [14], UNetFormer [29], MFTransNet [15], FTransUNet [16], and RS3Mamba [50], most of which were specifically designed for remote sensing tasks. In our experiments, ABCNet, PSPNet, MAResU-Net, UNetFormer, and RS3Mamba utilized only the optical images, to highlight the impact of DSM data and demonstrate the advantages of multimodal methods over single-modal ones. The other methods are state-of-the-art multimodal models based on different network architectures, including CNN and Transformer. The comparative quantitative results are presented in Table II and Table III.
@@ IV-C1 Performance Comparison on the Vaihingen dataset
As presented in Table II, the proposed MANet demonstrated substantial improvements in terms of OA, mF1 and mIoU metrics compared to existing segmentation methods. These results confirmed that our MANet can effectively leverage the extensive general knowledge embedded in SAM. In particular, MANet outperformed state-of-the-art models across three specific classes, namely Building, Tree and Impervious surface. In terms of the overall performance, the proposed MANet achieved an OA of $92.97\%$, an mF1 score of $91.71\%$ and a mIoU of $85.03\%$, reflecting gains of $0.57\%$, $0.50\%$ and $0.80\%$ over the second best method FTransUNet, respectively. Moreover, each of the three MANet backbone variants offers unique advantages. Even the smallest variant, ViT-B, is comparable to most methods, further validating that our multimodal fine-tuning approach can efficiently utilize general knowledge from SAM to assist with the semantic segmentation of multimodal remote sensing data. This result demonstrated the practical value of the proposed MANet and the MMAdapter in guiding the introduction of foundation models, like SAM, into multimodal remote sensing tasks.
@@
Fig. 7 presents a visual comparison of the results produced by various methods and the proposed MANet. MANet demonstrates superior performance in generating sharper and more precise boundaries for ground objects, such as trees, cars, and buildings, resulting in clearer separations. This also helps preserve the integrity of the ground objects. Overall, the visualizations produced by MANet exhibit a cleaner and more organized appearance. We attribute these improvements primarily to SAM’s powerful feature extraction capabilities. By applying the multimodal fine-tuning mechanism, SAM’s ability to segment every natural element is effectively extended to ground objects.
@@
TABLE IV: Quantitative results on the Vaihingen dataset with different modalities and fine-tuning mechanism. The best results are in Bold. (%).
@@
NIIRG The standard Adapter 96.29 93.09 80.15 89.08 92.59 92.02 90.94 83.69
NIIRG + DSM The standard Adapter 98.74 91.98 82.17 88.59 93.25 92.73 91.23 84.25
NIIRG + DSM MANet with MMAdapter 98.84 93.17 81.16 89.23 93.39 92.93 91.51 84.72
IV-C2 Performance Comparison on the Potsdam dataset
Experiments on the Potsdam dataset yielded results consistent with those from the Vaihingen dataset. As shown in Table III, the corresponding OA, mF1 and mIoU values were $91.71\%$, $92.70\%$, $86.69\%$ respectively, which corresponds to increases of $0.37\%$, $0.29\%$ and $0.49\%$, respectively, over FTransUNet. Notably, significant gains were observed for Building, Tree and Low Vegetation compared to other state-of-the-art methods. The MANet with smaller backbones also shows superior performance. This flexibility allows MANet to balance hardware requirements and performance needs across various application scenarios.
Fig. 8 shows a visualization example from Potsdam, where we observed more defined boundaries and intact object representations. These visual improvements are consistent with the mF1 and mIoU metrics shown in TableIII. Undoubtedly, it further validates the practical applicability of the proposed MANet and MMAdapter.
In additional, it is also observed that MANet is hard to achieve the best results simultaneously in identifying Tree and Low Vegetation. This challenge arises because both categories are characterized by irregular boundaries. Furthermore, their similarity, as well as the staggered or overlapping distribution, make them difficult to be distinguished from each other. Combining SAM with more specialized design to tackle the identification of these challenging categories presents an interesting future direction.
IV-D Modality Analysis
@@
To illustrate the necessity of multimodal fine-tuning, we conducted a modality analysis, with the results presented in Table IV. In the first experiment, we only used single-modality data and applied the standard Adapter mechanism to fine-tune the SAM’s image encoder. This experiment highlights the importance and need for multimodal information. In the second experiment, the SAM’s image encoder retained the standard Adapters but excluded the proposed MMAdapter. Therefore, this experiment could still extract remote sensing multimodal features independently, but it lacked crucial information fusion during the encoding process.
@@
Inspection of Table IV reveals the significant effectiveness of incorporating multimodal information. The improvement is particularly pronounced in the categories of Building and Impervious surface, which tend to have significant and stable surface elevation information. Additionally, this enhancement strengthens the model’s ability to distinguish other categories. Overall, the integration of multimodal information provides comprehensive benefits across the board. The introduction of the MMAdapter enables more effective utilization of DSM information, significantly enhancing the model’s ability to extract and fuse multimodal information. As a result, the semantic segmentation performance is further improved.
@@ IV-E Ablation Study
The proposed MANet consists of two core components: the SAM’s image encoder with MMAdapter and the DFM. To validate their effectiveness, ablation experiments were conducted by systematically removing specific components. As shown in Table V, two ablation experiments were designed. In the first experiment, the DFM was removed from MANet, leading to a lack of deep analysis and integration of high-level abstract remote sensing semantic features. The second experiment follows the setup of the second experiments in Table IV.
Figure 9: Three groups of heatmaps. (a) NIIRG images, (b) DSM, (c) heatmaps from NIIRG images and (d) heatmaps from DSM both generated by the original SAM, (e) heatmaps from NIIRG images and (f) heatmaps from DSM both generated by the proposed MMAdapter. The high-value areas in the heatmaps indicate objects identified as buildings by the methods. The effectiveness of our MMAdapter can be clearly observable.
@@
It is important to note that removing all Adapters from the SAM’s image encoder severely degrades the model’s performance, as confirmed in Fig. 9, which also illustrates the effectiveness of the multimodal fine-tuning mechanism. Inspection of Fig. 9(c) and (e) reveal that SAM, without fine-tuning, cannot extract meaningful features from remote sensing data, rendering it unsuitable for semantic segmentation tasks. However, inspection of Fig. 9(d) and (f) shows that, after fine-tuning with the MMAdapter, the heatmaps change dramatically. Furthermore, Fig. 9(e) clearly demonstrates that SAM, despite being trained on RGB optical images, is also effective when applied to non-optical DSM data. It is observed that DSM can effectively provide supplementary information. Therefore, the fine-tuned SAM’s image encoder is capable of recognizing and segmenting remote sensing objects effectively in multimodal tasks.
@@
TABLE V: Ablation study of the proposed MANet. The best results are in Bold.
MMAdapter DFM OA(%) mF1(%) mIoU(%)
92.73 91.23 84.25
92.80 91.30 84.35
92.93 91.51 84.72
@@
Inspection of Table V indicates that both the MMAdapter and DFM are essential for enhancing the performance of the proposed MANet. Specifically, the MMAdapter facilitates continuous information fusion, allowing for extraction and fusion of multimodal information as the encoding depth increases. The DFM verifies the importance of high-level features in the semantic segmentation of remote sensing data. In this work, we primarily introduce a new framework for leveraging SAM, rather than emphasize the high-level feature fusion techniques. Replacing DFM with a more advanced fusion model is expected to result in further performance improvement.

### v2

@@ IV-B Implementation details
All experiments were conducted using PyTorch on a single NVIDIA GeForce RTX 3090 GPU with 24GB of RAM. The stochastic gradient descent algorithm was used to train all the models under consideration with a learning rate of $0.01$, momentum of $0.9$, weight decay of $0.0005$, and a batch size of $10$. The batch size is reduced to $4$ when ViT-H is applied to meet the memory limit. All models were trained for a total of $50$ epochs, with each epoch comprising $1000$ batches. Basic data augmentation techniques, including random rotation and flipping, are applied after sample collection with the sliding window of size $256\times 256$. For MMAdapter, the down projection ratio is set to $0.25$. For MMLoRA, we follow the low-rank value in [hu2021lora] as $4$. More details can be found in https://github.com/sstary/SSRS.
To assess the semantic segmentation performance on multimodal remote sensing data, we use overall accuracy (OA), mean F1 score (mF1), and mean intersection over union (mIoU). These standard metrics enable a fair comparison between the proposed MFNet and other state-of-the-art methods. Specifically, OA evaluates both foreground classes and the background class, while mF1 and mIoU are calculated for the five foreground classes.
@@ IV-C Performance Comparison
We benchmarked the proposed MFNet against fifteen state-of-the-art methods, including PSPNet [PSPNet], MAResU-Net [MAResUNet], vFuseNet [vfusenet], FuseNet [FuseNet], ESANet [ESANet], SA-GATE [SAGATE], CMGFNet [CMGFNet], TransUNet [Transunet], CMFNet [ma2022crossmodal], UNetFormer [wang2022unetformer], MFTransNet [he2023mftransnet], FTransUNet [ma2024ftransunet], RS3Mamba [ma2024rs], FTransDeepLab [FTransDeepLab], and MultiSenseSeg [MultiSenseSeg], most of which were specifically designed for remote sensing tasks. In our experiments, PSPNet, MAResU-Net, UNetFormer, and RS3Mamba utilized only the optical images, to highlight the impact of DSM data and demonstrate the advantages of multimodal methods over single-modal ones. The other methods are state-of-the-art multimodal models based on different network architectures, including CNN, Transformer and Mamba. Taking into account the three different backbones provided by SAM, and the two multimodal fine-tuning architectures proposed in this work, we present six sets of experimental results for each dataset. The comparative quantitative results are presented in Table I and Table II.
@@ IV-C1 Performance Comparison on the Vaihingen dataset
As presented in Table I, the proposed MFNet demonstrated substantial improvements in terms of OA, mF1 and mIoU metrics compared to existing segmentation methods. These results confirmed that our MFNet can effectively leverage the extensive general knowledge embedded in SAM. In particular, MFNet outperformed state-of-the-art models across four specific classes, namely Building, Tree, Low Vegetation and Impervious surface. In terms of the overall performance, the proposed MFNet (MMAdapter) with ViT-H achieved an OA of $92.97\%$, an mF1 score of $91.71\%$ and a mIoU of $85.03\%$, reflecting gains of $0.24\%$, $0.29\%$ and $0.50\%$ over the second best method MultiSenseSeg, respectively. Moreover, each of the three MFNet backbone variants offers unique advantages. Even the smallest variant, ViT-B, is comparable to most methods, further validating that our multimodal fine-tuning framework can efficiently utilize general knowledge from SAM to assist with the semantic segmentation of multimodal remote sensing data. This result demonstrated the practical value of the proposed MFNet and MMAdapter or MMLoRA in guiding the introduction of foundation models, like SAM, into multimodal remote sensing tasks. On the other hand, we observed that MFNet based on MMLoRA performed less effectively than the MMAdapter-based MFNet, which will be analyzed in Sec. IV-F.
@@
Fig. 8 presents a visual comparison of the results produced by various methods and the best MFNet (MMAdapter) with ViT-H. MFNet demonstrates superior performance in generating sharper and more precise boundaries for ground objects, such as trees, cars, and buildings, resulting in clearer separations. This also helps preserve the integrity of the ground objects. Overall, the visualizations produced by MFNet exhibit a cleaner and more organized appearance. We attribute these improvements primarily to SAM’s powerful feature extraction capabilities. By applying the multimodal fine-tuning mechanism, SAM’s ability to segment every natural element is effectively extended to ground objects.
TABLE III: Quantitative results on the MMHunan dataset. The best results are in Bold. The second best results are underlined. (%).
Method Backbone OA mF1 mIoU
Cropland Forest Grassland Wetland Water Unused Land Built-up Area Total
FuseNet [FuseNet] VGG16 76.10 89.21 37.11 8.37 70.12 72.75 16.99 76.35 54.80 42.30
CMFNet [ma2022crossmodal] VGG16 83.83 82.27 43.88 39.57 70.14 81.72 23.66 77.41 59.63 46.52
MFTransNet [he2023mftransnet] ResNet34 78.99 91.30 33.80 26.80 77.38 78.29 27.17 80.07 61.45 48.25
FTransUNet [ma2024ftransunet] R50-ViT-B 78.75 90.54 32.13 27.51 76.64 75.59 48.63 79.47 61.95 48.78
CMGFNet [CMGFNet] ResNet34 82.08 87.90 37.50 26.48 79.82 74.64 41.15 79.85 62.69 49.44
FTransDeepLab [FTransDeepLab] ResNet101 79.39 88.89 35.71 30.88 83.95 78.14 32.60 80.62 62.51 49.66
MultiSenseSeg [MultiSenseSeg] Segformer-B2 78.03 90.93 40.47 38.16 80.19 81.03 38.14 80.51 63.74 50.76
MFNet (MMLoRA) ViT-B 76.83 90.69 24.16 22.03 80.17 78.12 29.66 79.35 58.79 46.54
ViT-L 79.39 88.89 35.71 30.88 83.95 78.14 32.60 80.62 62.51 49.66
ViT-H 76.42 90.65 38.08 20.78 74.48 77.93 38.69 78.87 60.38 47.63
MFNet (MMAdapter) ViT-B 74.81 91.47 27.59 25.00 86.47 78.92 37.46 80.68 62.10 49.20
ViT-L 79.19 89.52 46.07 42.23 81.15 77.58 39.65 80.93 65.33 51.82
ViT-H 79.66 90.06 42.61 38.81 80.31 78.92 40.04 81.07 64.13 51.08
Figure 10: Visualized comparisons on the MMHunan test set with the size of $256\times 256$. $1024\times 1024$. (a) IRRG images, (b) DSM, (c) Ground Truth, (d) CMFNet, (e) FTransUNet, (f) MFTransNet, (g) CMGFNet, (h) FTransDeepLab, (i) MultiSenseSeg, (j) The proposed MFNet. To highlight the differences, some purple boxes are added to all subfigures.
IV-C2 Performance Comparison on the Potsdam dataset
Experiments on the Potsdam dataset yielded results consistent with those from the Vaihingen dataset. As shown in Table II, the corresponding OA, mF1 and mIoU values of MFNet (MMAdapter) with ViT-H were $91.71\%$, $92.70\%$, $86.69\%$ respectively, which corresponds to increases of $0.37\%$, $0.29\%$ and $0.49\%$, respectively, over FTransUNet. Notably, significant gains were observed for Building, Tree, Low Vegetation and Impervious Surface compared to other state-of-the-art methods. The MFNet with smaller backbones also shows great performance. This flexibility allows MFNet to balance hardware requirements and performance needs across various application scenarios.
Fig. 9 shows a visualization example from Potsdam, where we observed more defined boundaries and intact object representations. These visual improvements are consistent with the mF1 and mIoU metrics shown in TableII. Undoubtedly, it further validates the practical applicability of the proposed MFNet and MMAdapter or MMLoRA.
Additionally, it is observed that MFNet struggles to achieve optimal performance in identifying both Tree and Low Vegetation simultaneously. This challenge arises because both categories are characterized by irregular boundaries.Furthermore, their similarity, combined with their staggered or overlapping distribution, makes it challenging to distinguish between them. Combining SAM with more specialized designs to tackle the identification of these challenging categories presents an interesting future direction.
IV-C3 Performance Comparison on the MMHunan dataset
The experimental results on the MMHunan dataset are presented in Table III. MFNet (MMAdapter) with a ViT-L backbone achieved OA, mF1, and mIoU scores of $80.93\%$, $65.33\%$, $51.82\%$, respectively, representing improvements of $0.42\%$, $1.59\%$ and $1.06\%$ over MultiSenseSeg. These results validate the consistent overall performance gains brought by our method. In addition, we observed an interesting phenomenon: across both fine-tuning strategies, ViT-H underperformed compared to ViT-L on this dataset. This suggests that in scenarios with small datasets, larger backbones may be more prone to overfitting, highlighting the importance of backbone selection for different remote sensing scenes. Fig. 10 shows a visualization example from MMHunan. In large-scale scenarios, the domain gap between remote sensing and natural images is more significant. However, our proposed approach successfully bridges this gap, allowing the vision foundation model’s general understanding to be effectively transferred to remote sensing tasks, resulting in consistent performance improvements.
@@ IV-C3 Performance Comparison on the MMHunan dataset
TABLE IV: Quantitative results on the Vaihingen dataset with different modalities and fine-tuning mechanisms. (%).
@@
NIIRG Without Adapter 94.64 89.47 71.71 76.83 89.51 88.01 85.34 75.11
NIIRG Standard LoRA 96.50 93.62 80.35 86.32 92.78 92.00 90.35 82.74
NIIRG + DSM Standard LoRA 97.26 92.61 81.58 86.53 91.58 91.86 89.90 82.06
NIIRG + DSM MFNet with MMLoRA 96.85 92.89 81.09 89.95 93.28 92.22 91.09 83.96
NIIRG Standard Adapter 96.29 93.09 80.15 89.08 92.59 92.02 90.94 83.69
NIIRG + DSM Standard Adapter 99.02 91.68 83.04 89.71 92.90 92.80 91.30 84.35
NIIRG + DSM MFNet with MMAdapter 98.84 93.17 81.16 89.23 93.39 92.93 91.51 84.72
IV-D Modality and Fine-tuning Analysis
@@
To illustrate the necessity of multimodal fine-tuning framework, we conducted the modality and fine-tuning analysis, with the results presented in Table IV. In the first experiment, we only used single-modality data and did not apply any fine-tuning mechanism, while in the second and fifth experiments, we applied standard Adapter/LoRA mechanisms to fine-tune the SAM’s image encoder. These experiments highlight the importance and need for multimodal information and fine-tuning mechanisms. In the third and sixth experiments, the SAM’s image encoder retained standard Adapter/LoRA but excluded the proposed MMAdapter/MMLoR. Therefore, these experiments could still extract remote sensing multimodal features independently, but they lacked crucial information fusion during the encoding process. The fourth and seventh experiments contain multimodal information and the proposed MMAdapter/MMLoR.
Inspection of Table IV first highlights the need for fine-tuning mechanisms. Without Adapter or LoRA, SAM struggles to effectively extract remote sensing features, leading to a significant decline in performance. Furthermore, the results reveal the great effectiveness of incorporating multimodal information. The improvement is particularly pronounced in the categories of Building and Impervious surface, which tend to have significant and stable surface elevation information. After that, this enhancement strengthens the model’s ability to distinguish other categories. Additionally, we observe that the performance in the third experiment is lower than in the second. This is attributed to low-rank decomposition significantly reducing the dimensionality of task-specific information. Hereby, the heterogeneity between modalities complicates their fusion after the encoder. It highlights the challenge that improper fine-tuning poses in leveraging multimodal information. Our MMLoRA effectively addresses this challenge through a gradual feature fusion approach in the image encoder.
@@
Overall, the integration of multimodal information provides comprehensive benefits across the board. The introduction of the MMAdapter and MMLoRA enables more effective utilization of DSM information, significantly enhancing the model’s ability to extract and fuse multimodal information. As a result, the semantic segmentation performance is further improved.
Figure 11: Four groups of heatmaps. (a) NIIRG images, (b) DSM, (c) heatmaps from NIIRG images and (d) heatmaps from DSM both generated by the original SAM, (e) heatmaps from NIIRG images and (f) heatmaps from DSM both generated by the proposed MMAdapter, (g) heatmaps from NIIRG images and (h) heatmaps from DSM both generated by the proposed MMLoRA. The high-value areas in the heatmaps indicate objects identified as buildings by the methods. The effectiveness of our MMAdapter and MMLoRA can be clearly observable.
@@ IV-E1 Component Ablation
The proposed MFNet consists of two core components: the SAM’s image encoder with MMAdapter or MMLoRA and the DFM. To validate their effectiveness, ablation experiments were conducted by systematically removing specific components. As shown in Table V, two ablation experiments were designed. In the first experiment, the DFM was removed from MFNet, leading to a lack of deep analysis and integration of high-level abstract remote sensing semantic features. The second experiment follows the setup of the third and sixth experiments in Table IV.
@@
Before analyzing the ablation experiments, it is important to note that removing all Adapters or LoRAs from the SAM’s image encoder severely degrades the model’s performance, as confirmed in Fig. 11, which also illustrates the effectiveness of the multimodal fine-tuning mechanisms. Inspection of Fig. 11(c) and (d) reveals that SAM, without fine-tuning, cannot extract meaningful features from remote sensing data, rendering it unsuitable for semantic segmentation tasks. However, inspection of Fig. 11(d) and (f), or Fig. 11(g) and (h) shows that, after fine-tuning with the MMAdapter or MMLoRA, the heatmaps change dramatically. Furthermore, Fig. 11(f) and (h) clearly demonstrate that SAM, despite being trained on RGB optical images, is also effective when applied to non-optical DSM data. It is observed that DSM can effectively provide supplementary information. Therefore, the fine-tuned SAM’s image encoder is capable of recognizing and segmenting remote sensing objects effectively in multimodal tasks.
@@
TABLE V: Ablation study of the proposed MFNet. The best results are in Bold.
MMAdapter DFM OA(%) mF1(%) mIoU(%) MMLoRA DFM OA(%) mF1(%) mIoU(%)
92.73 91.23 84.25 92.02 90.64 83.24
92.80 91.30 84.35 91.86 89.90 82.06
92.93 91.51 84.72 92.22 91.09 83.96
@@
Inspection of Table V indicates that both the multimodal fine-tuning and DFM are essential for enhancing the performance of the proposed MFNet. Specifically, the MMAdapter and MMLoRA facilitate continuous information fusion, allowing for extraction and fusion of multimodal information as the encoding depth increases. The DFM verifies the importance of high-level features in the semantic segmentation of remote sensing data. In this work, we primarily introduce a new framework for leveraging SAM, rather than emphasize the high-level feature fusion techniques. Replacing DFM with a more advanced fusion model is expected to result in further performance improvement.
Figure 12: Relationship between training data volume and model performance during the training stage. As the model is unable to generate predictions in the absence of training data, we designed and carried out experiments with three levels of training data availability: 25%, 50%, and 75%, in addition to using the complete training set (100%).
IV-E2 Data Amount Ablation
To investigate the fine-tuning efficiency of SAM on remote sensing tasks, we conducted experiments using varying proportions of the training data to explore the relationship between training data volume and model performance. Specifically, we fine-tuned the model using only 25%, 50%, and 75% of the training set, while evaluating on the full test set. The results, shown in Fig. 12, reveal a phenomenon where data quantity plays a crucial role between 25% and 50%, yet the performance gains tend to saturate beyond the 50% threshold. This suggests that SAM is capable of rapidly acquiring task-specific knowledge through fine-tuning, making further increases in training data yield diminishing returns in downstream performance. This finding offers valuable insights into the data requirements for related tasks, providing guidance on the efficient use of training data.
TABLE VI: Model scale analysis measured by a $256\times 256$ image on a single NVIDIA GeForce RTX 3090 GPU. For different MFNet configurations, the parameter statistics are: the fine-tuning parameters in SAM’s image encoder + the parameters in DFM and the decoder. MIoU values are the results of the Vaihingen dataset. The best results are in Bold.
Method Parameter (M) Memory (MB) MIoU (%)
PSPNet [PSPNet] 46.72 3124 76.96
MAResU-Net [MAResUNet] 26.27 1908 79.89
UNetFormer [wang2022unetformer] 24.20 1980 81.97
RS3Mamba [ma2024rs] 43.32 1548 82.78
TransUNet [Transunet] 93.23 3028 78.26
FuseNet [FuseNet] 42.08 2284 78.71
vFuseNet [vfusenet] 44.17 2618 78.92
ESANet [ESANet] 34.03 1914 79.42
SA-GATE [SAGATE] 110.85 3174 81.27
CMFNet [ma2022crossmodal] 123.63 4058 81.44
MFTransUNet [he2023mftransnet] 43.77 1549 81.61
CMGFNet [CMGFNet] 64.20 2463 82.26
FTransUNet [ma2024ftransunet] 160.88 3463 84.23
FTransDeepLab [FTransDeepLab] 69.86 1624 83.87
MultiSenseSeg [MultiSenseSeg] 60.46 2264 84.53
MFNet (MMLoRA) (ViT-B) 1.03+6.22 1924 82.09
MFNet (MMLoRA) (ViT-L) 2.75+6.22 4158 83.96
MFNet (MMLoRA) (ViT-H) 4.59+6.22 6520 84.66
MFNet (MMAdapter) (ViT-B) 14.20+6.22 1872 83.24
MFNet (MMAdapter) (ViT-L) 50.45+6.22 $42$ 84.72
MFNet (MMAdapter) (ViT-H) 105.06+6.22 6854 85.03